OpenStudy (anonymous):

am i correct on this? I CHOSE B 4. What is the solution set to the equation 2a2 + 8a = 0? (Points : 1) {2, 4} {0, 4} {0, -4} {-4,- 4}

Directrix (directrix):

I did not get option B. Let me check my work.

OpenStudy (anonymous):

i tried factoring like in the lesson then i just filled them in, i may be missing something again

Directrix (directrix):

@IDAREYOU If you substitute 4 back into this --> 2a2 + 8a = 0 you will not get zero.

OpenStudy (anonymous):

i didnt get B either, but B is the second best answer!

OpenStudy (anonymous):

even with -4?

Directrix (directrix):

@IDAREYOU Will you consider posting your factoring?

Directrix (directrix):

If this option is B {0, 4}, where is the -4? :)

OpenStudy (anonymous):

2a^2 +8a = 0 8a = 2 x 4 right?

Directrix (directrix):

2a(a + 4) = 0 2a = 0 or a + 4 = 0 a = ? or a = ?

Directrix (directrix):

@IDAREYOU Do you follow the factoring I did? I used the distributive property and then the zero product property.

OpenStudy (anonymous):

@IDAREYOU the answer is C. because when you put in 0=a you get the rhs to be a zero and when you put in a=-4 you get rhs to be a zero hence the option is C

OpenStudy (anonymous):

0/ 2 = 0 and a + 4 = 0 = a = 0 - 4 = a = -4

OpenStudy (anonymous):

do you want more work?

Directrix (directrix):

@croslali Yes, we believe in work to support answers.

Directrix (directrix):

@IDAREYOU 2a(a + 4) = 0 2a = 0 or a + 4 = 0 a = 0 or a = -4 Yes, you completed that correctly. Suggestion: Be systematic with your work. It will pay off.

OpenStudy (anonymous):

so if i check it: 2(-4)^2 + 8(-4) = 0 2 x 16 = 32 8 x -4 = -32 32 - 32 = 0 i wasn't using the ()

Directrix (directrix):

That is half the checking.

OpenStudy (anonymous):

i jumped ahead of myself on this one

Directrix (directrix):

Right. There is no prize for finishing first so be systematic. "Fast is fine but accuracy is everything." ~ Wyatt Earp

OpenStudy (anonymous):

2(0)^2 + 8(0) 2 x 0 + 8 x 0 0 + 0 = 0

Directrix (directrix):

Yes. That's the way to do these.

OpenStudy (anonymous):

ok, so when factoring its just the GCF?

Directrix (directrix):

On that particular problem, yes. On something like: x^2 + 5x - 6 = 0, then that is not what I think of as a GCF problem.

OpenStudy (anonymous):

so on this my gcf would be 10 thus i would have x = 0 x+5 = 0? 5. What is the solution set to the equation x2 + 10x = 21? (Points : 1) {3, 7} {-21,- 31} {0, -10} {-3, -7}

OpenStudy (anonymous):

oh wait that should be brought over like they showed so i would subtract or add 21 to both sides

Directrix (directrix):

There is no GCF here that I see. This is factoring a trinomial. x2 + 10x = 21 x^2 + 10 x - 21 = 0

OpenStudy (anonymous):

do you have microsoft word so you can see how they are showing it in my lesson?

Directrix (directrix):

@IDAREYOU Check and see if you posted the problem correctly. None of the options fit this: x2 + 10x = 21

OpenStudy (anonymous):

its a -21

OpenStudy (anonymous):

x2 + 10x = -21

Directrix (directrix):

I can read Word documents.

Directrix (directrix):

x2 + 10x = -21 x2 + 10x + 21 = 0 (x + ?)*(x + ?) = 0

OpenStudy (anonymous):

(x + 10) (x + 21) or (x + 5) (x + 7)

Directrix (directrix):

(x + 10) (x + 21) = x^2 + 31x + 210 so that's not it. (x + 5) (x + 7) = x^2 + 12x + 35 so that's not it.

Directrix (directrix):

x2 + 10x + 21 = 0 What are two numbers that multiply to 21 AND also add to 10 ?

Directrix (directrix):

@IDAREYOU

OpenStudy (anonymous):

3 X 7 = 21 7 + 3 = `10

Directrix (directrix):

x2 + 10x + 21 = x^2 + 7x + 3x + 21 = x* (x + 7 ) + 3* (x + 7) = (x + 7) * (x + 3)

Directrix (directrix):

x2 + 10x + 21 = 0 (x + 7) * (x + 3) = 0 (x + 7) = 0 OR (x + 3) = 0 x = ? or x = ?

OpenStudy (anonymous):

x = 0 - 7 x = 0 - 3

Directrix (directrix):

Yes. And, then x = -7, x = -3

OpenStudy (anonymous):

x = -7 x = 3 [-7,-3}

OpenStudy (anonymous):

i have 2 more quizzes to take before i win this bet with my dad

Directrix (directrix):

x = -7 x = 3 [-7,+3} --> You had a bad sign on the 3.

OpenStudy (anonymous):

um wait 3, - 7

Directrix (directrix):

I don't understand how this came about: x = -7 x = 3 [-7,-3}

Directrix (directrix):

x = -7, x = -3

OpenStudy (anonymous):

3 x 3 = 9 + 10 x 3 = 30 9 + 30 = 39 not 21 thats why i had a negative

OpenStudy (anonymous):

but -7 x -7 = 49 + 10 x -7 =-70 49 + -70 = -21 like in the original problem

Directrix (directrix):

two numbers that multiply to 21 AND also add to 10 are 7 and 3. They are not the solutions to the equation. They are numbers that figure into the factoring through which you get the solutions to the equation.

OpenStudy (anonymous):

yes but after i solve for zero x + 7 = 0 and x + 3 = 0

OpenStudy (anonymous):

i mean solve for x

Directrix (directrix):

Then, you get -7 and -3 as solutions to the equation.

OpenStudy (anonymous):

and that where i got (-7,-3) :)

Directrix (directrix):

Right. Seriously, you should work to get a grip on this factoring. It will follow you in problems for the next 3 years at least.

OpenStudy (anonymous):

Thank you

Directrix (directrix):

You are welcome.